OpenStudy (anonymous):

which number is between 0.5 and 5/8

OpenStudy (anonymous):

.5 is 1/2 5/8 is a little more then 1/2

OpenStudy (anonymous):

.625 i believe is 5/8 so a number between .5 and .625

OpenStudy (anonymous):

my answer choices are 1/4 , 8/28, 9/16, and 19/28

OpenStudy (igreen):

Simplify them all to decimals

OpenStudy (anonymous):

1/4 is way less then 1/2 it would have to be 2/4 8/28 is way less then 1/2

OpenStudy (igreen):

5/8 = 0.625 So we want a number between 0.5 and 0.625

OpenStudy (igreen):

Simplify the answer choices to decimals, can you do that? @Caleb_Smith102

OpenStudy (anonymous):

already got all dat :(

OpenStudy (anonymous):

find something that is more then 1/2 but less then 1/3 xD

OpenStudy (anonymous):

no its all in fractions.

OpenStudy (anonymous):

in ur answer choices

OpenStudy (igreen):

Yes, but if you plug them in google or a calculator you will get a decimal.

OpenStudy (jordan123321):

you can change a fraction in to a decimal

OpenStudy (igreen):

1/4 = ? 8/28 = ? 9/16 = ? 19/28 = ?

OpenStudy (anonymous):

it all needs a common denominator

OpenStudy (jordan123321):

igreen is telling you the right thing

OpenStudy (jordan123321):

you can change a fraction in to a decmial

OpenStudy (anonymous):

find the dicimal of the question being asked

OpenStudy (anonymous):

so something between .5 and .625 or whatever

OpenStudy (anonymous):

its really simple but u make it complicated

OpenStudy (anonymous):

something above 1/4

OpenStudy (anonymous):

thank yall. I got it right

OpenStudy (anonymous):

xD

OpenStudy (igreen):

What did you pick?

OpenStudy (anonymous):

something i was right about xD

OpenStudy (anonymous):

1/4 , 8/28, 9/16, and 19/28

OpenStudy (jordan123321):

@Caleb_Smith102 what did you pick

OpenStudy (anonymous):

something less then 19/28 something more then 8/28 and 1/4

OpenStudy (anonymous):

thats what he picked

OpenStudy (anonymous):

this is really simple...

OpenStudy (anonymous):

yeh no kidden

OpenStudy (anonymous):

its c...

OpenStudy (texaschic101):

to do problems such as these, there are 2 ways : you can either turn them all into fractions with common denominators..and compare them, or you can turn them all into decimals and compare them. Either way, you should arrive at the same answer

OpenStudy (anonymous):

or u could just be smart and look at it and pay attention in class